OpenStudy (anonymous):

Use the elimination method to solve the following system of equations. x + 3y – z = 2 x – 2y + 3z = 7 x + 2y – 5z = –21 (2, 3, 5) (–2, 3, 5) (2, –3, 5) (2, 3, –5) Use the elimination method to solve the following system of equations. x + 3y – z = 2 x – 2y + 3z = 7 x + 2y – 5z = –21 (2, 3, 5) (–2, 3, 5) (2, –3, 5) (2, 3, –5) @Mathematics

jhonyy9 (jhonyy9):

assume second with 3rd and will get 2x-2z= -14 and multiply first by 2 and second by 3 and assum and will get 5x +7z = 25 so than we have 2x-2z = -14 multiply by -1 than will be -2x +2z = 14 5x-2z = 25 -2x+2z=14 5x-2z=25 ------------- assume and will get 3x = 39 x= 39/3 = 13 x=13 so than 5(13) -2z=25 so 65 -2z = 25 so 65-25 =2z so 40=2z so z= 40/2 = 20 z=20 than x+3y-z=2 so 13 +3y -20 =2 so 3y-7=2 so 3y=9 so y=9/3 =3 than we will get x=13 y=3 z=20